OpenStudy (anonymous):

PLEASE HELP!!!! URGENT!!! Explain why the following statement is NOT true: "The day before a big election, a pre-election survey showed that Candidate A got 52%, and Candidate B got 48%. The margin of error is 3%. If nothing changes, Candidate A will win. " Thanks! =)

OpenStudy (anonymous):

If the margin of error is 3%, then A could be as low as 52 -3 = 49%. or as high as 52 +3 = 55% and, also, B could be as low as 48 -3 = 45% or as high as 48 + 3 = 51%.

OpenStudy (anonymous):

a=0.03 is really a small ratio, so we assume its nearly impossible for a point to locate in a<=0.03 that area, then we say A will win

OpenStudy (anonymous):

So, it is possible that A is at 49% and B is at 51%, so B could be winning if nothing changes.

OpenStudy (anonymous):

You can show this by finding the highest and lowest the percents could be using the 3% margin of error. For example, if you calculate the president with the 48% as the highest it can be and the lowest it can be using the 3% and this percent is still less than the other president when the same process is applied to that president, the statement is true.

OpenStudy (anonymous):

Wow! Thanks SO much, everybody! :) I so appreciate this!

OpenStudy (anonymous):

I hope you were able to find the answer. :)